OpenStudy (anonymous):

Someone please help me!!! graph y(less than or equal to) -2x+3

jimthompson5910 (jim_thompson5910):

step 1) graph the line y = -2x + 3 step 2) make it a solid line (we're including the boundary) step 3) shade below the solid line (there's a "less than" in the inequality sign)

OpenStudy (anonymous):

How do I graph it though ? that's what I don't get .

jimthompson5910 (jim_thompson5910):

to graph any linear equation, you just need 2 points

jimthompson5910 (jim_thompson5910):

to find a point, you plug in an x value to get a corresponding y value

jimthompson5910 (jim_thompson5910):

for instance, if x = 0, y = -2x + 3 y = -2*0 + 3 y = 0 + 3 y = 3 giving us one point (0,3)

jimthompson5910 (jim_thompson5910):

repeat that for a different x value (say x = 1) to get another point